OpenStudy (anonymous):

Solve x2 + 2x + 9 = 0. x equals negative 2 plus or minus 4 I square root of 2 x equals negative 2 plus or minus 2 I square root of 2 x equals negative 1 plus or minus 4 I square root of 2 x equals negative 1 plus or minus 2 I square root of 2

OpenStudy (anonymous):

\[x^2+2x+9=0\\ x^2+2x=-9\] is a start then what is half of 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

ok and \{1^2=1\) so next step is \[(x+1)^2=-9+1\\ (x+1)^2=-8\]

OpenStudy (anonymous):

you good from there? take the square root, get \[x+1=\pm\sqrt{-8}\] and so on

OpenStudy (anonymous):

i really have no clue how to do any of this